OpenStudy (anonymous):

1. What is the slope of the line that passes through the given points? (2, 12) and (6, 11) (1 point) –1 over 4 1 over 4 4 –4

OpenStudy (cggurumanjunath):

m=(y2-y1)/(x2-x1)

OpenStudy (cggurumanjunath):

(x1,y1)=(2,12) (x2,y2)=(6,11)

OpenStudy (anonymous):

that would be -1 over 4 using the slope formula (*the one given by guy above me :P *) that is m=(y2-y1)/(x2-x1) slope (m) = (11-12)/(6-2) = -1/4 the order in which you take the points does not matter!! that is wether you take (2,12) or (6,11) as (X1,Y1) does not matter youll get the same answer :P Good Luck!! =D

OpenStudy (anonymous):

Thank you guys!!

OpenStudy (cggurumanjunath):

:)